OpenStudy (anonymous):

int (e^ (root x)/(2root x) dx

OpenStudy (amistre64):

\[\int_{} \frac{e^{\sqrt{x}}}{2\sqrt{x}}dx\] this?

OpenStudy (amistre64):

\[e^\sqrt{x}+C\]